OpenStudy (anonymous):

What price do farmers get for their watermelon crops? In the third week of July, a random sample of 42 farming regions gave a sample mean of x = $6.88 per 100 pounds of watermelon. Assume that σ is known to be $1.82 per 100 pounds. (a) Find a 90% confidence interval for the population mean price (per 100 pounds) that farmers in this region get for their watermelon crop. What is the margin of error? (Round your answers to two decimal places.)

OpenStudy (anonymous):

eww stats. Sorry mate. I don't know anything about confidence intervals.

OpenStudy (anonymous):

this is the only thing about stats i dont get lol

OpenStudy (zarkon):

use z interval for the mean...you should have the formula

OpenStudy (zarkon):

\[\bar{x}\pm z_{\alpha/2}\frac{\sigma}{\sqrt{n}}\]

OpenStudy (anonymous):

so if i look up the z interval for 6.88 i can then use that in the formula

OpenStudy (zarkon):

you are doing a 90% CI...you need to use this number to find \[z_{\alpha/2}\]

OpenStudy (anonymous):

which in non-mathamatical terms means what :)

OpenStudy (zarkon):

\[1-\alpha=.9\] so \[\alpha=.1\] \[\frac{\alpha}{2}=.05\] find \[z_{.05}\]

OpenStudy (anonymous):

ah thnx

OpenStudy (zarkon):

Find the number \[z\] such that \[P(Z>z)=.05\]